Find all numbers for which the rational experssion is undefined.
t^3-7t/t^2-81
Thanks for the help

Answer by RAY100(1637)

(t^3 -7t) / (t^2-81)
.
undefined if denominator is zero
.
t^2-81 = (t+9)(t-9) =0,,,t=+9,,-9